OpenStudy (anonymous):

Your friend hands you a graph of the engine power statistics of a race car. He says, “I know this graph is f(x) = –3(x + 4)^6 – 8 but I can’t remember how it is related to the graph of x^4.” Explain to your friend how the graph f(x) is a translation of the graph x^4. I get translations, but I can't seem to identify this one.

OpenStudy (anonymous):

i hope there is a typo there

OpenStudy (anonymous):

\[ f(x) = –3(x + 4)^6 – 8\] should maybe be \[ f(x) = –3(x + 4)^4 – 8\]

OpenStudy (anonymous):

it is shifted a few ways

OpenStudy (anonymous):

4 units to the left flipped about the \(x\) axis because of the \(-3\) then down \(8\) units because of the \(-8\) also narrower because of the \(x^6\)

OpenStudy (anonymous):

it is the minus sign that flips it over the \(x\) axis

OpenStudy (anonymous):

compare say \(y=x^2\) to \(y=-x^2\)

OpenStudy (anonymous):

and \(x^6\) is narrow than \(x^4\) because raising a number to the power of 6 makes it a lot bigger than raising it to the power of 4 (assuming it is larger than 1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i mean yes the \(x+4\) shifts it 4 units left

OpenStudy (anonymous):

sure what is the other one ?

OpenStudy (anonymous):

they really stretch to make these word problem, don't they? "break even" they tell you just means \(p(x)=0\) so you have to find the zeros of this thing your first answer is correct. lets find the zeros

OpenStudy (anonymous):

yes stay graphical the zeros are not rational

OpenStudy (anonymous):

well really it would not have domain \(\mathbb{R}\) but rather \((0,\infty)\) since you do not make a negative amount of cans

OpenStudy (anonymous):

yes